Q. ∫_(-∞)^∞▒e^(〖-x〗^2 ) dx=?
a. π/2
b. π
c. √π
d. 2√π
1
Expert's answer
2019-07-09T11:51:49-0400

The given definite integration is a special integration ,

"erf\left( x \right) = \frac{1}{{\sqrt \pi }}\int_{ - x}^x {{e^{ - {t^2}}}dt} = \frac{2}{{\sqrt \pi }}\int_{ 0}^x {{e^{ - {t^2}}}dt}\\
\int_{ - x}^x {{e^{ - {t^2}}}dt} = \frac{{\sqrt \pi }}{2}\int_{ - x}^x {\frac{{2{e^{ - {t^2}}}}}{{\sqrt \pi }}dt}"



Now, use this concept and our integration will be

"\int_{ - \infty }^\infty {{e^{ - {x^2}}}} dx = \frac{{\sqrt \pi }}{2}\int_{ - \infty }^\infty {\frac{{2{e^{ - {x^2}}}}}{{\sqrt \pi }}dx} \\
 = \frac{{\sqrt \pi }}{2} \cdot \left[ {erf(x)} \right]_{ - \infty }^\infty \\
 = \frac{{\sqrt \pi }}{2} \cdot 2\\
 = \sqrt \pi"

Therefore, the option C is the right answer.